# Characterization of fractional Chern insulator quasiparticles in moiré transition metal dichalcogenides

### v1

@@ VI Conclusion
In this work, we have investigated the properties of quasiparticles for the $\nu_{h}=2/3$ FCI in tMoTe2 and for the $\nu=2/3$ FQH state in the LLL. We have proposed a tight-binding model for mobile FCI quasiparticles. We further pin quasiparticles by delta impurity potentials. We get clear signal of the nucleation of localized quasiparticles, as shown by the consistency of numerically obtained charge and braiding phase of quasiparticles with theoretical values. The spatial extent of a localized quasiparticle is also estimated by comparing the accumulated excess charge with the FQH counterpart. Trial wavefunctions of FCI quasiparticles are variationally determined using the nearly ideal quantum geometry of the tMoTe2 top valence band.
There are several possible future directions based on this work. First, our numerical characterization of FQH/FCI quasiparticles still suffers from non-negligible finite-size effects. Even within the single-band/LLL projection, the localized quasiparticle does not fully develop, as reflected by the oscillating excess charge and residual deviation of the braiding phase from the theoretical value. The $\nu=2/3$ quasihole is even much bigger, thus we cannot extract useful information about it by ED. We have also made tentative investigations of band mixing effects on FCI quasiparticles in tMoTe2 (see Appendix D), but the results are limited to smaller systems and much less conclusive. More advanced numerical techniques are hence necessary for solving these problems. If the recent applications of neural-network algorithms to FQH and FCI systems [104, 105, 106, 107] can reach larger systems than ED, they might be helpful to characterize the quasiparticle/quasihole more precisely, clarify their response to band mixing in moiré materials, and even study non-Abelian excitations. Second, we only considered the short-range delta impurity potential in this work. It would also be useful to study more realistic long-range impurity potentials, such as the Gaussian potential and the Coulomb potential of a point charge above the sample [85]. Finally, we focused on the anyons of stable $\nu=2/3$ Laughlin FCIs. As the non-uniform quantum geometry is a distinctive feature of moiré bands compared to Landau levels, it is interesting to systematically study the quantum geometric effect on the interaction between mobile FCI anyons and the spatial structure of localized FCI anyons. For the latter, a density wave halo around anyons is predicted to appear when FCIs compete with charge density waves [108].
@@ Appendix B Diagonalize the full Hamiltonian of tMoTe2
In Fig. 10, we compare the ground states obtained by this alternative method with those in Sec. IV for one and two localized quasiparticles. With the increasing of impurity potential strength $U_{0}$, the overlap grows quickly. For the single-quasiparticle case, the overlap has reached $\approx 94\%$ when $U_{0}=1\ {\rm meV}$. However, $U_{0}=4\ {\rm meV}$ is needed for the two-quasiparticle case if we want to reach the same high overlap, probably because the two quasiparticles are not well separated from each other in the finite system.
We hence estimate the impurity potential strength required to pin an isolate $\nu_{h}=2/3$ FCI quasiparticle as $\sim 1\ {\rm meV}$. This strength is still much smaller than the band gap of the top valence band in the single valley, which is $\approx 8.6\ {\rm meV}$. In particular, we also notice that the overlaps are still growing even when $U_{0}$ has exceeded the gap separating the quasiparticle subspace and higher excited levels (Fig. 4), indicating that the quasiparticle states cannot be mixed with other states by strong delta impurities. This also happens for the $\nu=1/3$ model Laughlin quasiholes, which are always the zero-energy ground states of the first-order Haldane pseudopotential combined with the delta pinning potential irrespective of the potential strength.
@@ Appendix D Two-band simulations
Figure 12: (a) Density profile around a single quasiparticle excitation in the $\nu_{h}=2/3$ tMoTe2 FCI with $N=11,N_{s}=16$. We set $N_{\rm b2}=N$, so hole’s occupation in the second valence band is unrestricted. The impurity potential (indicated by the black cross) is located at $R_{M}^{M}$. (b) Excess charge in the presence of a single-quasiparticle excitation in the $\nu_{h}=2/3$ tMoTe2 FCI with $N=13,N_{s}=19$. Here we compare the data obtained with different $N_{\rm b2}$. The horizontal line indicates $Q=e/3$, which is expected for a Laughlin-type quasiparticle.
In Fig. 12(a), we display the profile of a single quasiparticle for $N=11,N_{s}=16,N_{\rm b2}=N$, with the impurity potential located at the $R_{M}^{M}$ point. This is the largest system we can reach under two-band projection if the hole’s occupation in the second valence band is unrestricted. Clearly the quasiparticle is not well developed for this system size. Compared with Fig. 7, the density discrepancy still strongly oscillates in the entire sample. In Fig. 12(b), we gradually increase $N_{\rm b2}$ in a larger system size $N=11,N_{s}=19$ to demonstrate how the localized quasiparticle is affected by the band mixing. In the limit of $N_{\rm b2}=0$, the excess charge as a function of distance $r$ from the impurity potential is very similar to that shown in Fig. 7(d). With the increasing of $N_{\rm b2}$, the first peak of $Q(r)$ is gradually suppressed, such that $Q(r)$ does not oscillate around $e/3$ at the available values of $r$. We notice that the curve does not change much from $N_{\rm b2}=3$ to $N_{\rm b2}=4$, so probably $N_{\rm b2}=4$ is already close to the case of $N_{\rm b2}=N$.
@@
The appreciable suppression of the first peak of $Q(r)$ was also observed using the delta impurity potential for the $\nu=1/3$ Laughlin state of electrons in the second Landau level [85]. So it may be useful to compare the top and the second tMoTe2 valence bands with conventional Landau levels. In the aspect of single-particle quantum geometry, a convenient indicator of their similarity is

### v3

@@ VI Conclusion
In this work, we have investigated the properties of quasiparticles for the $\nu_{h}=2/3$ FCI in tMoTe2 and for the $\nu=2/3$ FQH state in the LLL. We have proposed a tight-binding model for mobile FCI quasiparticles. We further pin quasiparticles by delta impurity potentials. We get clear signal of the nucleation of localized quasiparticles, as shown by the consistency of numerically obtained charge and braiding phase of quasiparticles with theoretical values. The spatial extent of a localized FQH quasiparticle is estimated by either using the density profile or the braiding phase. The size of a FCI quasiparticle can then be estimated by mapping to its FQH counterpart via an effective lattice magnetic length. Trial wavefunctions of FCI quasiparticles are variationally determined using the nearly ideal quantum geometry of the tMoTe2 top valence band. The effective interaction between two $e/3$ FCI quasiparticles shows a crossover from repulsion to attraction when the gate screening in the Coulomb potential is enhanced, which may provide a way to form $2e/3$ anyons.
There are several possible future directions based on this work. First, our numerical characterization of FQH/FCI quasiparticles still suffers from non-negligible finite-size effects. Even within the single-band/LLL projection, the localized quasiparticle does not fully develop, as reflected by the oscillating excess charge and residual deviation of the braiding phase from the theoretical value. The $\nu=2/3$ quasihole is even much bigger, thus we cannot extract useful information about it by ED (see Appendix E). We have also made tentative investigations of band mixing effects on FCI quasiparticles in tMoTe2 (see Appendix D), but the results are limited to smaller systems and much less conclusive. More advanced numerical techniques are hence necessary for solving these problems. If the recent applications of neural-network algorithms to FQH and FCI systems [111, 112, 113, 114] can reach larger systems than ED, they might be helpful to characterize the quasiparticle/quasihole more precisely, clarify their response to band mixing in moiré materials, and even study non-Abelian excitations. Second, we only considered the short-range delta impurity potential in this work. It would also be useful to study more realistic long-range impurity potentials, such as the Gaussian potential and the Coulomb potential of a point charge above the sample [90]. Finally, we focused on the anyons of stable $\nu=2/3$ Laughlin FCIs. As the non-uniform quantum geometry is a distinctive feature of moiré bands compared to Landau levels, it is interesting to systematically study the quantum geometric effect on the interaction between mobile FCI anyons and the spatial structure of localized FCI anyons. For the latter, a density wave halo around anyons is predicted to appear when FCIs compete with charge density waves [115].
@@ Appendix B Diagonalize the full Hamiltonian of tMoTe2
In Fig. 10, we compare the three ground states obtained by this alternative method with those in Sec. IV for one and two localized quasiparticles. With the increasing of impurity potential strength $U_{0}$, the overlap grows quickly. For the single-quasiparticle case, the overlap has reached $\approx 94\%$ when $U_{0}=1\ {\rm meV}$. However, $U_{0}=4\ {\rm meV}$ is needed for the two-quasiparticle case if we want to reach the same high overlap, probably because the two quasiparticles are not well separated from each other in the finite system.
We hence estimate the impurity potential strength required to pin an isolate $\nu_{h}=2/3$ FCI quasiparticle as $\sim 1\ {\rm meV}$. This strength is still much smaller than the band gap of the top valence band in the single valley, which is $\approx 8.6\ {\rm meV}$. In particular, we also notice that the overlaps are still growing even when $U_{0}$ has exceeded the gap separating the quasiparticle subspace and higher excited levels (Fig. 4), indicating that the quasiparticle states cannot be mixed with other states by strong delta impurities. This also happens for the $\nu=1/3$ model Laughlin quasiholes in Eq. (32), which remain the zero-energy ground states of the first-order Haldane pseudopotential combined with the delta pinning potential irrespective of the potential strength. Therefore, we expect that the quasiparticle state in tMoTe2 is indeed close to the model Laughlin quasiparticle subject to density modulation, as shown in Eq. (35).
@@ Appendix C Dependence on dielectric constant
Figure 12: (a) Density profile around a single localized quasiparticle excitation in the $\nu_{h}=2/3$ tMoTe2 FCI with $N=11,N_{s}=16$. We set $N_{\rm b2}=N$, so hole’s occupation in the second valence band is unrestricted. The impurity potential (indicated by the black cross) is located at $R_{M}^{M}$. (b) Excess charge in the presence of a single localized quasiparticle excitation in the $\nu_{h}=2/3$ tMoTe2 FCI with $N=13,N_{s}=19$. Here we compare the data obtained with different $N_{\rm b2}$. The horizontal line indicates $Q=e/3$, which is expected for a Laughlin-type quasiparticle.
@@ Appendix D Two-band simulations
In Fig. 12(a), we display the profile of a single quasiparticle for $N=11,N_{s}=16,N_{\rm b2}=N$, with the impurity potential located at the $R_{M}^{M}$ point. This is the largest system we can reach under two-band projection if the hole’s occupation in the second valence band is unrestricted. The density has been averaged over the three nearly degenerate states with a single localized quasiparticle. Clearly the quasiparticle is not well developed for this system size. Compared with Fig. 7, the density discrepancy still strongly oscillates in the entire sample. In Fig. 12(b), we gradually increase $N_{\rm b2}$ in a larger system size $N=11,N_{s}=19$ to demonstrate how the localized quasiparticle is affected by the band mixing. In the limit of $N_{\rm b2}=0$, the excess charge as a function of distance $r$ from the impurity potential is very similar to that shown in Fig. 7(d). With the increasing of $N_{\rm b2}$, the first peak of $Q(r)$ is gradually suppressed, such that $Q(r)$ does not oscillate around $e/3$ at the available values of $r$. We notice that the curve does not change much from $N_{\rm b2}=3$ to $N_{\rm b2}=4$, so probably $N_{\rm b2}=4$ is already close to the case of $N_{\rm b2}=N$.
@@
The appreciable suppression of the first peak of $Q(r)$ was also observed using the delta impurity potential for the $\nu=1/3$ Laughlin state of electrons in the second Landau level [116, 90]. So it may be useful to compare the top and the second tMoTe2 valence bands with conventional Landau levels. In the aspect of single-particle quantum geometry, a convenient indicator of their similarity is
@@
Appendix E Localized $\nu=2/3$ quasiholes
Figure 14: (a), (b) The density distribution for a localized single-quasihole excitation created in (a) the $\nu=2/3$ FQH state on a square torus with $N=23,N_{\phi}=35$ and (b) the $\nu_{h}=2/3$ FCI in tMoTe2 with $N=23,N_{s}=35$. In both cases the impurity potential is located at the center of the sample, which coincides with the $\mathcal{R}_{\text{Br}}$ position in tMoTe2. (c) Excess charges around the impurity potential. In the FCI case we consider different positions of the impurity potential, as indicated in the legend. The horizontal reference line indicates Q=e/3Q=-e/3, which is expected for a Laughlin-type quasihole. We use the screened Coulomb interaction for both FQH and FCI systems.
In the main text, we have characterized localized $\nu=2/3$ quasiparticles for both the FQH and FCI cases. For completeness, we now present the results of a single localized $\nu=2/3$ quasihole. The calculations are similar to those in Secs. II.2 and IV.3, with the same model details. We use the LLL projection and the single-band projection for the FQH and FCI system, respectively. The subspace of a single mobile $\nu=2/3$ quasihole is created when the system size satisfies $N_{\phi/s}=\frac{1}{2}(3N+1)$. We then diagonalize a repulsive delta impurity potential to pin the quasihole. Three nearly degenerate ground states are obtained.
The density profiles in the LLL and tMoTe2 are displayed in Figs. 14(a) and 14(b), respectively, which have been averaged over the three nearly degenerate ground states. In both cases we indeed observe density minima near the impurity potential, which is expected for a localized quasihole. However, as shown in Figs. 14(c), the excess charge in the quasihole case is much messier than that in the quasiparticle case [Fig. 7(d)]. In particular, the excess charge does not start regular oscillations around e/3-e/3 at relatively small $r$. It is first stuck between about 0.2e-0.2e and $0.05e$ until $r\sim 4\ell_{B/0}$. Even after that we cannot see a clear tendency of convergence to e/3-e/3. Moreover, $Q(r)$ in tMoTe2 now shows strong dependence on the position of the impurity potential. The data indicate that the spatial extent of a localized $\nu=2/3$ quasihole is larger than that of a localized $\nu=2/3$ quasiparticle, so that we cannot characterize it well in finite systems reached by ED. In the context of FQH states, it has been known that the localized $\nu=1/3$ quasiparticle (equivalently, the $\nu=2/3$ quasihole) has a bigger size because of its complicated internal structure [117, 118].